# Clarification and Coinduction of Tambara Functors

### v1

@@ 5.7. Nullstellensatzian clarified Tambara functors
$\mathrm{Hom}_{\mathcal{C}}(x,-):\mathcal{C}\rightarrow\mathsf{Set}$
@@
For example, in rings, a $k$-algebra is compact if and only if it is finitely presented. In fact, this remains true in Tambara functors: [SSW25, Proposition 4.3] asserts that, if $k$ is a Tambara functor, then a $k$-algebra $R$ is compact (in the category of $k$-algebras) if and only if it is the coequalizer of a diagram $k[x_{G/H_{i}}]_{i\in I}\rightrightarrows k[x_{G/H_{j}}]_{j\in J}$. In this case we again call $R$ finitely presented.
@@
The idea is that Nullstellensatzian objects are abstractions of algebraically closed fields in a certain sense. Hilbert’s Nullstellensatz implies that a ring is Nullstellensatzian if and only if it is an algebraically closed field. One might therefore hope that Nullstellensatzian objects in equivariant algebra are related to $G$-actions on algebraically closed fields. One might hope that this relationship would have implications for the inverse Galois problem, as there is usually a plentiful supply of Nullstellensatzian objects in any reasonable category of “things that look like rings”. The author acknowledges David Chan for suggesting that there could be a relationship of this form.
@@
$\mathrm{Hom}(\mathcal{A}[x_{G/G}],\mathcal{A}[x_{G/G}])\cong\mathcal{A}[x_{G/G}](G/G)\mathrm{.}$
@@
$G/G\leftarrow G/e\rightarrow G/e\rightarrow G/e\mathrm{.}$
@@
Finally, it seems like the problem of analyzing Nullstellensatzian bi-incomplete $\Lambda$-clarified $G$-Tambara functors should be both interesting and difficult. Most of the results of this subsection should be true in this greater generality, so in particular it would be interesting to know what differences appear among Nullstellensatzian objects for different choices of $\Lambda$ and of compatible pairs of transfer systems.

### v2

@@ 5.2. Nullstellensatzian clarified Tambara functors
$\mathrm{Hom}_{\mathcal{C}}(x,-):\mathcal{C}\rightarrow\mathrm{Set}$
@@
For example, in rings, a $k$-algebra is compact if and only if it is finitely presented. In fact, this remains true in Tambara functors: [SSW25, Proposition 4.3] asserts that, if $k$ is a Tambara functor, then a $k$-algebra $R$ is compact (in the category of $k$-algebras) if and only if it is the coequalizer of a diagram $k[x_{G/H_{i}}]_{i\in I}\rightrightarrows k[x_{G/H_{j}}]_{j\in J}$ with $I$ and $J$ finite. In this case we again call $R$ finitely presented.
@@
The idea is that Nullstellensatzian objects are abstractions of algebraically closed fields in a certain sense. Hilbert’s Nullstellensatz implies that a ring is Nullstellensatzian if and only if it is an algebraically closed field. One might therefore hope that Nullstellensatzian objects in equivariant algebra are related to $G$-actions on algebraically closed fields. One might hope that this relationship would have implications for the inverse Galois problem, as there is usually a plentiful supply of Nullstellensatzian objects in any reasonable category of “things that look like rings”. The author acknowledges that the idea that there could be a relationship of this form is due to David Chan.
@@
$\mathrm{Hom}(\mathcal{A}[x_{G/G}],\mathcal{A}[x_{G/G}])\cong\mathcal{A}[x_{G/G}](G/G).$
@@
$G/G\leftarrow G/e\rightarrow G/e\rightarrow G/e.$